Find the perimeter of a polygon with vertices F(3,0), A(5,4), I(0,7), T(-5,4), and H(-3,0).
Answer by Boreal(15235)

5 distance formulas
FA is sqrt (2^2+4^2) or sqrt (20)
AI is sqrt(25+9)=sqrt(34)
IT is sqrt (34)
TH is sqrt (20)
HF is 6
this is 6 +2sqrt(20) + 2 sqrt(34)
=6+4 sqrt(5)+4 sqrt(17) units